# From MTEB to MTOB: Retrieval-Augmented Classification for Descriptive Grammars

### v1

@@ 6 Discussion
Expanding the MTOB approach could significantly benefit from standardizing descriptive grammars of various languages into a uniform format, leveraging databases such as Grambank or WALS. However, we can state that in the non-contaminated environment demonstrated in this work the descriptive formal linguistic texts still pose a significant challenge. While generally machine reading can be perceived as a “solved task”, the results on a lot of fine-grained linguistic features show that descriptive grammars remain a non-saturated material showcasing LLMs’ weak spots.
@@ 9 Ethical Statement
In pursuing the advancement of machine translation and natural language processing through the use of descriptive grammars, we must address critical ethical concerns, particularly relating to copyright. Descriptive grammars often come with copyright licenses which can pose significant barriers to their extensive utilization in large-scale machine reading projects.
To ensure compliance with copyright laws and respect for intellectual property rights, we publish the first benchmark fully, since publishing individual paragraphs for research purposes falls under fair use. Regarding the second benchmark, we publish the list of grammars and the list of page numbers where the relevant information for each feature can be found.
@@
Appendix A Benchmark for Rerankers: Examples for Relevance Categories
@@
Appendix B Benchmark for Rerankers: Details
@@
(Campbell, 2017) 34 7 0 2 0 7 0,68 0,18 0,14
(Newman, 2002) 15 30 3 1 1 0 0,30 0,68 0,02
(Georg, 2007) 36 13 0 0 1 0 0,72 0,26 0,02
(Forker, 2013) 9 14 10 6 4 7 0,18 0,60 0,22
(Ford, 1998) 28 19 0 0 0 3 0,56 0,38 0,06
(Tsunoda, 2012) 8 23 7 5 3 4 0,16 0,70 0,14
(Morgan, 1991) 0 12 10 11 4 13 0,00 0,66 0,34
(Dunn, 1979) 45 0 1 2 0 2 0,90 0,06 0,04
(Elliott, 2021) 5 18 2 4 3 18 0,10 0,48 0,42
(Sakel, 2004) 11 19 5 6 4 5 0,22 0,60 0,18
(Hellwig, 2019) 24 10 1 0 7 8 0,48 0,22 0,30
(Wegener, 2012) 17 16 7 3 1 6 0,34 0,52 0,14
(Olawsky, 2006) 6 17 3 2 8 14 0,12 0,44 0,44
(Weir, 1986) 34 4 2 2 5 3 0,68 0,16 0,16
Total 272 202 51 44 41 90 0,3886 0,4243 0,1871
@@
Appendix C Benchmark for Rerankers: Dynamics for NDCG@k
@@
Appendix D Benchmark for Rerankers: Wikipedia Summary vs. Term Only
@@
Appendix E Benchmark for RAG: Baseline Prompts
@@
Separate words (i. e. particles that are not phonologically bound to other words) are separated from other words by spaces.

### v2

@@ 6 Discussion
Expanding the MTOB approach could significantly benefit from standardizing descriptive grammars of various languages into a uniform format, leveraging databases such as Grambank or WALS. However, we can state that in the non-contaminated environment demonstrated in this work the descriptive linguistic texts still pose a significant challenge. While generally machine reading can be perceived as a “solved task”, the results on linguistic features show that descriptive grammars remain a non-saturated material showcasing LLMs’ weak spots.
@@ 8 Limitations
Furthermore, due to high computational costs of running RAG on GPT-4o with extensive excerpts from grammars, we limited our experiments (excluding the baseline, which was executed 10 times) to running every RAG configuration only once.
@@ 9 Ethical Statement
In pursuing the advancement of machine translation and natural language processing through the use of descriptive grammars, we must address critical ethical concerns, particularly relating to copyright. Most grammars in the benchmark have copyright licenses which prohibit their reproduction fully.
The benchmark for rerankers and the benchmark for RAG differ in the following way: in the benchmark for rerankers, each individual paragraph is annotated according to relevance to WALS 81A (to be used for retrieval assessment), whereas the benchmark for RAG does not have annotations for individual paragraphs. Therefore, we release our codebase and materials under MIT and publish the first benchmark fully, since it can be considered as a derivative work, and its publication for research purposes falls under fair use.
To ensure compliance with copyright laws and respect for intellectual property rights, we do not publish the second benchmark fully; we only publish the list of grammars and the list of page numbers where the relevant information for each feature can be found.
@@
Appendix A Rerankers: HuggingFace
  1. 1.
    bge-en-icl (Xiao et al., 2023): https://huggingface.co/BAAI/bge-en-icl
  2. 2.
    stella_en_1.5B_v5: https://huggingface.co/dunzhang/stella_en_1.5B_v5
  3. 3.
    NV-Retriever-v1 (de Souza et al., 2024): https://huggingface.co/nvidia/NV-Retriever-v1
  4. 4.
    gte-Qwen2-7B-instruct (Li et al., 2023): https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct
  5. 5.
    Linq-Embed-Mistral (Kim et al., 2024): https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral
  6. 6.
    SFR-Embedding-2_R (Meng* et al., 2024): https://huggingface.co/Salesforce/SFR-Embedding-2_R
  7. 7.
    SFR-Embedding-Mistral (Meng et al., 2024): https://huggingface.co/Salesforce/SFR-Embedding-Mistral
Appendix B Benchmark for Rerankers: Examples for Relevance Categories
@@
Appendix C Benchmark for Rerankers: Details
@@
(Campbell, 2017) 34 7 0 2 0 7 0.68 0.18 0.14
(Newman, 2002) 15 30 3 1 1 0 0.30 0.68 0.02
(Georg, 2007) 36 13 0 0 1 0 0.72 0.26 0.02
(Forker, 2013) 9 14 10 6 4 7 0.18 0.60 0.22
(Ford, 1998) 28 19 0 0 0 3 0.56 0.38 0.06
(Tsunoda, 2012) 8 23 7 5 3 4 0.16 0.70 0.14
(Morgan, 1991) 0 12 10 11 4 13 0.00 0.66 0.34
(Dunn, 1979) 45 0 1 2 0 2 0.90 0.06 0.04
(Elliott, 2021) 5 18 2 4 3 18 0.10 0.48 0.42
(Sakel, 2004) 11 19 5 6 4 5 0.22 0.60 0.18
(Hellwig, 2019) 24 10 1 0 7 8 0.48 0.22 0.30
(Wegener, 2012) 17 16 7 3 1 6 0.34 0.52 0.14
(Olawsky, 2006) 6 17 3 2 8 14 0.12 0.44 0.44
(Weir, 1986) 34 4 2 2 5 3 0.68 0.16 0.16
Total 272 202 51 44 41 90 0.3886 0.4243 0.1871
@@
Appendix D Benchmark for Rerankers: Dynamics for NDCG@k
@@
Appendix E Benchmark for Rerankers: Wikipedia Summary vs. Term Only
@@
Appendix F Benchmark for RAG: Prompts
@@
Separate words (i.e., particles that are not phonologically bound to other words) are separated from other words by spaces.
@@
Appendix G Benchmark for RAG: Prompt Details
G.1 Wikipedia Summaries
Feature Wikipedia Title
WALS 81A Word order [^3] https://en.wikipedia.org/w/index.php?title=Word_order&oldid=1240489972
GB 107 Affirmation and negation [^4] https://en.wikipedia.org/w/index.php?title=Affirmation_and_negation&oldid=1202783032
WALS 49A Grammatical case [^5] https://en.wikipedia.org/w/index.php?title=Grammatical_case&oldid=1238822129
WALS 116A* Yes-no question [^6] https://en.wikipedia.org/w/index.php?title=Yes%E2%80%93no_question&oldid=1236424936
Table 7: The Wikipedia pages we chose to take their summaries as definitions for the typological features (versions as of September 3rd, 2024).
The fourth paragraph from the "Grammatical case" summary was not included in the prompts due to containing examples of languages and their numbers of cases. If it was included in the prompt, we would have essentially provided the RAG system with an answer to a part of the given task.
G.2 Base Prompts
Due to the multilabel classification on the linguistic domain already posing a potential significant challenge for the RAG pipeline, we simplified the annotation guidelines for the features pertaining to interrogative particles compared to the ones presented in Grambank.
Firstly, in our version of the feature, the particles do not necessarily have to be dedicated to marking polar questions; the main purpose of the particle may be marking content questions, as long as using it to mark polar questions is possible as well. The motivation for this adjustment is the fact that the retriever and reranker in the RAG pipeline will retrieve information related to polar questions only based on the query, and it is irrational to expect information on polar questions to appear in the retrieved paragraphs.
Secondly, we deemed it unnecessary for a clause-medial particle to have the middle of the clause as its most common placement; a particle is considered clause-medial if it can appear in the clause-medial position in at least one example.
Another simplification is omission of two other features related to polar questions in Grambank: GB286, GB297. V-not-V constructions and double-marking with both a particle and verbal morphology occur in our benchmark in isolated examples only.
Despite the simplifications, we impose the requirement on assigning clitics to the particle category if they can attach to anything in their vicinity, and to the verb morphology category if they can be phonologically bound to the verb only, in order for the task to test the capabilities of the RAG pipeline in regard to distinguishing morphology from syntax. GB291 (marking polar questions by means of tone) has been retained for a similar purpose: distinguishing morphology from prosody.
G.3 Modifications for Chain-of-Thought
We modified each chapter of the WALS and Grambank for the Chain-of-Thought prompts in the following way according to our specifications: We eliminated all mentions and examples of languages that already exist in the respective benchmarks. For instance, (1) has been modified to obtain (2) due to Oneida being present in the Number of Cases benchmark:
(1) This excludes, for example, the "locative" suffixes in Oneida (Iroquoian; Ontario) from being counted as case, since they can derive body-part nouns which may occur in all semantically permitted syntactic positions (i.e. not only as locational adverbials). (WALS, Chapter 49A: Number of Cases)
(2) For example, "locative" suffixes that can derive body-part nouns which may occur in all semantically permitted syntactic positions (i.e. not only as locational adverbials), are not counted as case.
Additionally, we removed mentions of the WALS map, since each prompt for the RAG pipeline is formulated as determining a feature value instead of placing it on a map. We eliminated duplicated information (i.e., left only one mention about disregarding leading polar questions across all chapters in GB257 – GB 291); changed the guidelines according to our simplified definitions of clause-initial/clause-medial/clause-final particles; and included the explanation of when clitics are regarded as particles and when they are considered verbal morphology.
Appendix H Benchmark for RAG: Distribution of Values
Word Order (WALS 81A)
No dominant order 40
No mention 12
OSV 1
OVS 1
SOV 59
SVO 25
VOS 4
VSO 6
Total 148
Number of Cases (WALS 49A)
No case marking 55
Excl. borderline case marking 21
2 cases 11
3 cases 3
4 cases 5
5 cases 6
6-7 cases 22
8-9 cases 13
10 or more cases 12
Total 148
Standard Negation (GB 107)
0 65
1 83
Total 148
Polar Questions (WALS 116A*)
Clause-initial particle
0 20
1 128
Clause-medial particle
0 132
1 16
Clause-final particle
0 100
1 48
Verb morphology
0 125
1 23
Interrogative word order
0 144
1 4
Interrogative intonation only
0 96
1 52
Tone
0 141
1 7
Total 148
Table 8: Distribution of values in each feature (total = 148). The imbalance of classes should not be mitigated, because the benchmark, serving as a representative language sample, is intended to reflect the natural distribution of values found in languages around the world. If the grammar does not mention the feature, the following values are assigned to it: No mention for WALS 81A; No case marking for WALS 49A; 0 for GB 107; 0 for each of the sub-features for WALS 116A*. We make the assumption that a lack of mention in the descriptive material most likely indicates a lack of the phenomenon in the language (apart from the Word Order feature, since one of the seven logically possible categories must inevitably be present).
Appendix I Benchmark for Rerankers: Ablation
Additionally to the main experiment, we conducted an ablation study: how would the RAG pipeline perform if it had access to a near-optimal retriever instead of the combination of BM25 and rerankers, i.e., how much of the error is attributable to the generation component of Retrieval-Augmented Generation?
In order to conduct this ablation study, we annotated the pages in the grammar documents where the relevant information for the feature is contained (i.e., sufficient information for a linguist to unambiguously infer the value of the feature). Since this approach is only applicable to grammars that actually contain the relevant information, we ran the ablation on 136 grammars for WALS 81A, 146 grammars for GB 107, 121 grammars for WALS 116A, and 140 grammars for WALS 49A (out of 148), the same subsets as for the baseline runs.
We used the manually retrieved pages as the retrieved component in RAG, resulting in two columns on the right in Table 9, and recalculated the metrics for the four RAG configurations on the same subsets as used for the ablation. The baseline metrics are identical to the ones in Table 4 due to identical subsets of grammars.
The results indicate that for features that have a high baseline performance (most likely due to having an extreme class imbalance), e.g, Interrogative Word Order and Tone, involving a human to retrieve the paragraphs may harm the classification quality.
For WALS 49A (the feature hypothesized to be the most difficult one to classify) it is the Chain-of-Thought component that makes the difference between harming the classification performance and improving it; however, the F1 scores for this feature remain low, not exceeding 0.56.
F1 average Baseline Wiki
Wiki
+CoT
Re-
ranker
Rer.
+CoT
Human
Human
+CoT
micro 0.5551 ± 0.0359 0.7353 0.7426 0.7500 0.7794 0.7868 0.7868
macro 0.2812 ± 0.0276 0.7060 0.5866 0.5880 0.6000 0.5198 0.4949
WALS 81A weighted 0.5328 ± 0.0337 0.7398 0.7499 0.7391 0.7721 0.7838 0.7939
micro 0.5747 ± 0.0350 0.6644 0.6644 0.6849 0.7192 0.8151 0.8425
macro 0.5740 ± 0.0354 0.5679 0.5814 0.6100 0.6644 0.8038 0.8304
GB 107 weighted 0.5724 ± 0.0361 0.5958 0.6069 0.6334 0.6829 0.8102 0.8366
micro 0.3986 ± 0.0238 0.5000 0.5214 0.5357 0.5000 0.5000 0.5429
macro 0.2216 ± 0.0158 0.4655 0.4461 0.4299 0.4031 0.4126 0.4737
WALS 49A weighted 0.3453 ± 0.0237 0.5088 0.5341 0.5409 0.5023 0.5082 0.5579
micro 0.4504 ± 0.0396 0.8347 0.8926 0.8430 0.8678 0.9174 0.9256
macro 0.4240 ± 0.0456 0.8346 0.8924 0.8429 0.8677 0.9169 0.9251
interrog. intonation only weighted 0.4068 ± 0.0484 0.8352 0.8929 0.8433 0.8681 0.9178 0.9260
micro 0.9653 ± 0.0076 0.9917 0.9835 0.9835 0.9752 0.9917 0.9835
macro 0.6641 ± 0.0460 0.9423 0.8957 0.8707 0.8269 0.9264 0.8707
interrog. word order weighted 0.9611 ± 0.0064 0.9922 0.9851 0.9835 0.9765 0.9912 0.9835
micro 0.7264 ± 0.0402 0.9008 0.9091 0.9421 0.9256 0.9587 0.9339
macro 0.5446 ± 0.0502 0.8127 0.8385 0.8972 0.8679 0.9266 0.8889
clause- initial particle weighted 0.7371 ± 0.0337 0.8987 0.9100 0.9427 0.9263 0.9591 0.9362
micro 0.4909 ± 0.0409 0.7190 0.7851 0.7025 0.7851 0.8017 0.8017
macro 0.4808 ± 0.0415 0.7190 0.7826 0.7025 0.7826 0.8000 0.7977
clause- final particle weighted 0.4661 ± 0.0428 0.7195 0.7874 0.7030 0.7874 0.8038 0.8035
micro 0.6760 ± 0.0367 0.8430 0.8843 0.8430 0.8595 0.9091 0.9008
macro 0.5011 ± 0.0544 0.7961 0.8374 0.7961 0.8049 0.8671 0.8606
clause- medial particle weighted 0.6644 ± 0.0355 0.8502 0.8857 0.8502 0.8621 0.9085 0.9021
micro 0.7678 ± 0.0185 0.7686 0.8099 0.8017 0.8182 0.8760 0.8843
macro 0.4533 ± 0.0311 0.7026 0.7361 0.7279 0.7506 0.8188 0.8235
interrog. verb morphology weighted 0.7102 ± 0.0160 0.7894 0.8226 0.8157 0.8311 0.8819 0.8877
micro 0.9273 ± 0.0128 0.9174 0.9504 0.9587 0.9669 0.9504 0.9669
macro 0.5147 ± 0.0573 0.7500 0.8199 0.8572 0.8484 0.8199 0.8662
tone weighted 0.9104 ± 0.0121 0.9309 0.9555 0.9637 0.9669 0.9555 0.9689
Table 9: F1 scores for the ablation experiments (on the right) and the four initial RAG configurations + baseline (on the left) on the same subset of grammars as the ablation experiments. The best F1 value in each row in the main table and the best F1 value for the better of the two ablation experiments are shown in bold. Each of the two ablation experiments was executed only once. Blue cells in the ablation experiments indicate better performance, red cells indicate worse performance, and purple cells indicate identical performance compared to the the best non-human retriever in the same row.
Most features demonstrate an increase in F1 scores when compared to non-human retrievers.
The greatest improvement in quality was achieved on the balanced binary value, GB 107: +0.166 macro F1 compared to the best RAG configuration.
Overall, a conclusion can be made that while the task of machine reading on the linguistic domain is still not fully resolved, the task of information retrieval within the linguistic domain is similarly important.